OpenStudy (anonymous):

Which equations in point-slope form are equations of the line that passes through the point (−7, 5) and (−4, 1)? Choose exactly two answers that are correct.

OpenStudy (anonymous):

\[A. y -7=-\frac{ 3 }{ 4 }\left( x- 5\right)\] \[B. y -5=-\frac{ 4 }{ 3 }\left( x +7 \right)\] \[C. y +4=-\frac{ 3 }{ 4 }\left( x-1 \right)\] \[D. y-1=-\frac{ 4 }{ 3 }\left( x+4 \right)\]